If city A is 450 miles east of city B, and car C leaves A for B at 50 miles an hour, and car D leaves City B for city A at 60 miles an hour, at what point does car C meet D?
Answer by stanbon(75887)

If city A is 450 miles east of city B, and car C leaves A for B at 50 miles an hour, and car D leaves City B for city A at 60 miles an hour, at what point does car C meet D?
-----------------------
A to B DATA:
rate = 50 mph ; time = x hours ; distance = 50x miles
-------------------------
B to A DATA:
rate = 60 mph ; time = x hours ; distance = 60x miles
-------------------------
Equation:
miles + miles = 450 miles
50x + 60x = 450
110x = 450
x = 4.09 hours
------------------
50x = 50*4.09 = 204.55 miles from A toward B
60x = 60*4.09 = 245.45 miles from B toward A
=================================================
Cheers,
Stan H.